# centertableaux

### v1

@@ 1.2. Representation-theoretic and combinatorial notations
for the positive root system, the root system, the positive coroot system, and coroot system, respectively, for the reductive Lie algebra $\mathfrak{g}$. Where we use a superscript * to denote the linear dual of a vector space. Let $Q_{\mathfrak{g}}\subseteq{{}^{a}\mathfrak{h}}^{*}$ and $Q^{\mathfrak{g}}\subseteq{{}^{a}\mathfrak{h}}^{*}$ denote the root lattice and the weight group of $\mathfrak{g}$, respectively. Let $W\subseteq{\mathrm{GL}}(^{a}\mathfrak{h})$ denote the abstract Weyl group. By the Harish-Chandra isomorphism, there is a 1-1 correspondence between $W$-orbits of $\nu\in{{}^{a}\mathfrak{h}^{*}}$ and algebraic characters $\chi_{\nu}:{\mathcal{Z}}(\mathfrak{g})\to{\mathbb{C}}$. We say that an ideal of ${\mathcal{U}}(\mathfrak{g})$ has infinitesimal character $\nu$ if it contains the kernel of $\chi_{\nu}$.
@@
Suppose $\nu\in{{}^{a}\mathfrak{h}^{*}}$ and ${\mathcal{O}}$ is a nilpotent orbit in $\mathfrak{g}^{*}$, denote by ${\mathrm{Irr}}_{\nu}(G;{\mathcal{O}})$ the subset of ${\mathrm{Irr}}(G)$ consisting of isomorphism classes of irreducible representations with infinitesimal character $\nu\in{{}^{a}\mathfrak{h}^{*}}$ and complex associated variety $\overline{{\mathcal{O}}}$ (the Zariski closure of ${\mathcal{O}}$). Then, we observe that for any ${\mathrm{Inn}}(\mathfrak{g})$-stable Zariski closed subset $S$ of ${\mathrm{Nil}}(\mathfrak{g}^{*})$
@@
However, in order to deal with the real unitary case, we need to introduce an additional setting where $\star=\widetilde{A}$, $G=\widetilde{U}(p,q)$ and $G_{{\mathbb{C}}}={\mathrm{GL}}_{n}({\mathbb{C}})$ ($n=p+q$). Here, $\widetilde{{\mathrm{U}}}(p,q)$ denotes the (linear) double cover of ${\mathrm{U}}(p,q)$, defined via the pullback diagram.
@@ 1.3. The main results
In the statement of the following theorems, we will use the usual identification ${}^{a}\mathfrak{h}^{*}={\mathbb{C}}^{n}\ \textrm{or}\ {\mathbb{C}}^{n}\times{\mathbb{C}}^{n}$ defined in Section 3.1, and identify nilpotent orbits in Lie algebras of type $A$ with partitions or Young diagrams in the usual way. For a partition $\iota$ denote by ${\mathcal{O}}_{\iota}$ the corresponding nilpotent orbit, and for a nilpotent orbit ${\mathcal{O}}$, denote by $\iota({\mathcal{O}})$ the corresponding partition. We also define the “row by row” union of Young diagrams $\mathop{\sqcup}\limits^{r}$ by
@@
where $k^{\prime}\geq k$ are positive integers. For example, $[5,3,1]\mathop{\sqcup}\limits^{r}[4,3,3,3]=[9,6,4,3]$
@@
    {ytableau}&$\ ~{}~{}\ and\ ~{}\ytableau~{}&~{}~{}.$
@@
    {ytableau}&={ytableau}$\ ~{}~{}\mathop{\sqcup}\limits^{r}~{}\ytableau~{},$
    {ytableau}&={ytableau}&$\ytableau~{}&~{}~{}~{}=~{}\ytableau~{}&~{}~{}\mathop{\sqcup}\limits^{r}~{}\ytableau~{}.$
    After enumerating all the assignments and paintings, we find that $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}_{1}))=4$ and $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}_{2}))=3$
@@
    (1.9) $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}))=\left\{\begin{aligned} &\sum_{\begin{subarray}{c}(\iota_{1},\cdots,\iota_{r})\in\mathrm{YD}_{e_{1}}\times\cdots\times\mathrm{YD}_{e_{r}}\\ \iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}=\iota({\mathcal{O}})\end{subarray}}\prod_{i=1}^{r}\sharp\left({\mathrm{Irr}}_{{\boldsymbol{\lambda}}_{i}}\left({\mathrm{GL}}_{\frac{e_{i}}{2}}({\mathbb{H}});{\mathcal{O}}_{\iota_{i}}\right)\right),&\textrm{if each $e_{i}$ is even}\\ &0,&\textrm{otherwise}.\end{aligned}\right.$
@@
    Suppose $G={\mathrm{GL}}_{4}({\mathbb{H}})$, $\nu=(1,1,1,2,2,2,3,3)$. Then, the associated variety of irreducible Casselman-Wallach representations with this infinitesimal character can only be the closure of the nilpotent orbit ${\mathcal{O}}$, whose Young diagram is
@@
    {ytableau}&$\ ~{}\ and\ ~{}\ytableau~{}&~{}~{}~{}\\ ~{}~{}~{}~{}\\ ~{}\\ ~{}.$
@@
    (1.10) $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}))=\left\{\begin{aligned} &0,&\textrm{if $\lambda_{1}-\lambda_{1}^{\prime}\notin{\mathbb{Z}}$},\\ &\sharp\left({\mathrm{A}}_{[d_{1},\cdots,d_{k}]}(\iota)\right)\cdot\sharp\left({\mathrm{A}}_{[d_{1}^{\prime},\cdots,d_{l}^{\prime}]}(\iota^{\prime})\right)\cdot\delta_{\iota,\iota^{\prime}},&\textrm{if $\lambda_{1}-\lambda_{1}^{\prime}\in{\mathbb{Z}}$}\end{aligned}\right.$
@@
    $\delta_{\iota,\iota^{\prime}}=\left\{\begin{aligned} &0,&\textrm{if $\iota\neq\iota^{\prime}$}\\ &1,&\textrm{if $\iota=\iota^{\prime}$}.\end{aligned}\right.$
@@ 2.1. Associated varieties and genuine representations
Following the notations introduced in Section 1, denote by $({\mathcal{U}}(\mathfrak{g})_{k})_{k=0,1,2,\cdots}$ the natural increasing filtration of ${\mathcal{U}}(\mathfrak{g})$, where ${\mathcal{U}}(\mathfrak{g})_{k}$ is the subspace of ${\mathcal{U}}(\mathfrak{g})$ spanned by elements of the form $X_{1}X_{2}\cdots X_{m}$ for $0\leq m\leq k$, with $X_{j}\in\mathfrak{g}$ for $1\leq j\leq m$. By the Poincaré-Birkhoff-Witt theorem, we can identify the associated ring
@@
is automatically a ${\mathcal{Z}}(\mathfrak{g})$-finite two-sided ideal, we define the complex associated variety $\mathrm{AV}(V)$ of $V$ as the associated variety of the ideal ${\mathrm{Ann}}(V)$, which is ${\mathrm{Inn}}(\mathfrak{g})$-stable and contained in ${\mathrm{Nil}}(\mathfrak{g}^{*})$. Suppose $S$ is a ${\mathrm{Inn}}(\mathfrak{g})$-stable Zariski closed subset of ${\mathrm{Nil}}(\mathfrak{g}^{*})$, denote by ${\mathrm{Rep}}_{S}(G)$ the category of Casselman-Wallach representations of $G$ whose complex associated variety is contained in $S$.
@@
$\psi:$ ${\mathrm{Irr}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q);{\mathcal{O}}\right)$$\longrightarrow$ $\quad{\mathrm{Irr}}_{\nu+\mu}\left({\mathrm{U}}(p,q);{\mathcal{O}}\right)$
@@ 2.2. Generalities on coherent families of Casselman-Wallach representations
It can be verified directly from the above assumption that the homomorphism $\iota$ induces a unique morphism between the abstract Cartan subalgebra of ${\mathrm{Lie}}(G)$ and that of ${\mathrm{Lie}}(G_{{\mathbb{C}}})$. The analytic weight lattice $Q_{an}$ of $G_{\mathbb{C}}$ is identified with a subgroup of ${{}^{a}\mathfrak{h}^{*}}$ via this morphism, and is denoted by $Q_{\iota}\subseteq{{}^{a}\mathfrak{h}^{*}}$. Moreover, $Q_{\iota}$ is $W$-stable and $Q_{\mathfrak{g}}\subseteq Q_{\iota}\subseteq Q^{\mathfrak{g}}$. In the rest of this section, we fix a $Q_{\iota}$-coset $\Lambda=\lambda+Q_{\iota}\subseteq{{}^{a}\mathfrak{h}^{*}}$. Let $W_{\Lambda}$ denote the stabilizer of $\Lambda$ in $W$. Specifically
@@ A ${\mathcal{K}}(G)$-valued $\Lambda$-coherent family is a map
We then define the notion of parameters for coherent continuation representations and compute them explicitly. Suppose $H$ is a Cartan subgroup of $G$, meaning that it is the centralizer of a Cartan subalgebra of $\mathrm{Lie}(G)$ in $G$. Since $G$ is in Harish-Chandra’s class, $H$ has a unique maximal compact subgroup $T$. Denote by $\Delta_{\mathfrak{h}}\subseteq\mathfrak{h}^{*}$ the root system of $\mathfrak{g}$. A root is called imaginary if $\check{\alpha}\in\mathfrak{t}$, where $\mathfrak{t}=\mathrm{Lie}(T)$, or equivalently if it takes purely imaginary values on $\mathfrak{h}^{*}$. Moreover, an imaginary root $\alpha$ is called compact imaginary if the corresponding root space $\mathfrak{g}_{\alpha}$ is contained in a complexified Lie algebra of a compact subgroup of $G$.
@@
  • Every irreducible Casselman-Wallach representation of $H$ has finite dimension.
@@
$\Gamma:$ $\Lambda$$\longrightarrow$ $\quad{\mathrm{Irr}}(H)$
@@
where $W(\Delta_{\mathrm{\mathfrak{h},im}})$ is the Weyl group for the imaginary root system $\Delta_{\mathfrak{h},\mathrm{im}}$. There is a group action of $W_{\mathfrak{h},\mathfrak{t}}$ on the set of positive systems of $\Delta_{\mathfrak{h},\mathrm{im}}$, with the subgroup $W(\Delta_{\mathfrak{h},\mathrm{im}})$ acts simply transitively. It is easy to verify that for any choice of positive systems $\Delta_{\mathfrak{h},\mathrm{im}}^{+}$ we obtain a decomposition:
@@
$\mathrm{sgn_{\gamma}}:$ $W_{\gamma}$$\longrightarrow$ $\quad{\mathbb{C}}^{\times}$
@@
2.3. Representation of Weyl groups
@@
Denote by $\mathrm{S}^{e}(V_{1})$ denote the $e$-th symmetric power of $V_{1}$. Suppose $\sigma_{1}\in{\mathrm{Irr}}(H_{1})$ is an irreducible representation that occurs with multiplicity $1$ in $\mathrm{S}^{e}(V_{1})$ and does not occur in $\mathrm{S}^{i}(V_{1})$ if $10\leq i\leq e-$ 1. We may regard $\sigma_{1}$ as a subspace of $\mathrm{S}^{e}(V)$ via the inclusion $\sigma_{1}\subseteq{\mathbb{C}}\otimes\mathrm{S}^{e}(V_{1})\subseteq\mathrm{S}^{e}(V)$, and consider the $H$-representation $\sigma$ of $\mathrm{S}^{e}(V)$ generated by $\sigma^{\prime}$. Then
@@
$j_{H_{1}}^{H}:$ ${\mathrm{Irr}}(H_{1},e)$$\longrightarrow$ $\quad{\mathrm{Irr}}(H,e)$
@@
Back to our setting of real reductive groups, $W\subseteq{\mathrm{GL}}({{}^{a}\mathfrak{h}})$ denote the abstract Weyl group. For every $\sigma\in{\mathrm{Irr}}(W)$, its fake degree is defined as
@@
ϕ::italic-$\phi:$ $\mathrm{YD}_{n}$$\longrightarrow$ $\quad{\mathrm{Irr}}(\mathrm{S}_{n})$
@@
When we identify nilpotent orbits with partitions of $n$, the bijection $\phi$ coincides with the Springer correspondence.

### v2

@@ 1.2. Representation-theoretic and combinatorial notations
for the positive root system, the root system, the positive coroot system, and coroot system, respectively, for the reductive Lie algebra $\mathfrak{g}$. Here we use a superscript * to denote the linear dual of a vector space. Let $Q_{\mathfrak{g}}\subseteq{{}^{a}\mathfrak{h}}^{*}$ and $Q^{\mathfrak{g}}\subseteq{{}^{a}\mathfrak{h}}^{*}$ denote the root lattice and the weight group of $\mathfrak{g}$, respectively. Let $W\subseteq{\mathrm{GL}}(^{a}\mathfrak{h})$ denote the abstract Weyl group. By the Harish-Chandra isomorphism, there is a 1-1 correspondence between $W$-orbits of $\nu\in{{}^{a}\mathfrak{h}^{*}}$ and algebraic characters $\chi_{\nu}:{\mathcal{Z}}(\mathfrak{g})\to{\mathbb{C}}$. We say that an ideal of ${\mathcal{U}}(\mathfrak{g})$ has infinitesimal character $\nu$ if it contains the kernel of $\chi_{\nu}$.
@@
Suppose $\nu\in{{}^{a}\mathfrak{h}^{*}}$ and ${\mathcal{O}}$ is a nilpotent orbit in $\mathfrak{g}^{*}$, denote by ${\mathrm{Irr}}_{\nu}(G;{\mathcal{O}})$ the subset of ${\mathrm{Irr}}(G)$ consisting of isomorphism classes of irreducible representations with infinitesimal character $\nu\in{{}^{a}\mathfrak{h}^{*}}$ and complex associated variety $\overline{{\mathcal{O}}}$ (the Zariski closure of ${\mathcal{O}}$). Then we observe that for any ${\mathrm{Inn}}(\mathfrak{g})$-stable Zariski closed subset $S$ of ${\mathrm{Nil}}(\mathfrak{g}^{*})$
@@
However, in order to deal with the real unitary case, we need to introduce an additional setting where $\star=\widetilde{A}$, $G=\widetilde{{\mathrm{U}}}(p,q)$ and $G_{{\mathbb{C}}}={\mathrm{GL}}_{n}({\mathbb{C}})$ ($n=p+q$). Here, $\widetilde{{\mathrm{U}}}(p,q)$ denotes the (linear) double cover of ${\mathrm{U}}(p,q)$, defined via the pullback diagram.
@@ 1.3. The main results
In the statement of the following theorems, we will use the usual identifications ${}^{a}\mathfrak{h}^{*}={\mathbb{C}}^{n}\ \textrm{or}\ {\mathbb{C}}^{n}\times{\mathbb{C}}^{n}$ defined in Section 3.1, and identify nilpotent orbits in Lie algebras of type $A$ with partitions or Young diagrams in the usual way. For a partition $\iota$ denote by ${\mathcal{O}}_{\iota}$ the corresponding nilpotent orbit, and for a nilpotent orbit ${\mathcal{O}}$, denote by $\iota({\mathcal{O}})$ the corresponding partition. We also define the “row by row” union of Young diagrams $\mathop{\sqcup}\limits^{r}$ by
@@
where $k^{\prime}\geq k$ are positive integers. For example, $[5,3,1]\mathop{\sqcup}\limits^{r}[4,3,3,3]=[9,6,4,3]$.
@@
    {ytableau}&$\ ~{}\quad and\quad\ytableau~{}&~{}~{}.$
@@
    {ytableau}&={ytableau}$\ ~{}\quad\mathop{\sqcup}\limits^{r}\quad\ytableau~{},$
    {ytableau}&={ytableau}&$\ytableau~{}&~{}~{}\quad=\quad\ytableau~{}&~{}\quad\mathop{\sqcup}\limits^{r}\quad\ytableau~{}.$
    After enumerating all the assignments and paintings, we find that $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}_{1}))=4$ and $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}_{2}))=3$.
@@
    (1.9) $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}))=\left\{\begin{aligned} &\sum_{\begin{subarray}{c}(\iota_{1},\cdots,\iota_{r})\in\mathrm{YD}_{e_{1}}\times\cdots\times\mathrm{YD}_{e_{r}}\\ \iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}=\iota({\mathcal{O}})\end{subarray}}\prod_{i=1}^{r}\sharp\left({\mathrm{Irr}}_{{\boldsymbol{\lambda}}_{i}}\left({\mathrm{GL}}_{\frac{e_{i}}{2}}({\mathbb{H}});{\mathcal{O}}_{\iota_{i}}\right)\right),&\textrm{if each $e_{i}$ is even;}\\ &0,&\textrm{otherwise}.\end{aligned}\right.$
@@
    Suppose $G={\mathrm{GL}}_{4}({\mathbb{H}})$, $\nu=(1,1,1,2,2,2,3,3)$. Then the associated variety of irreducible Casselman-Wallach representations with this infinitesimal character can only be the closure of the nilpotent orbit ${\mathcal{O}}$, whose Young diagram is
@@
    {ytableau}&$\ \quad and\quad\ytableau~{}&~{}~{}~{}\\ ~{}~{}~{}~{}\\ ~{}\\ ~{}.$
@@
    (1.10) $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}))=\left\{\begin{aligned} &0,&\textrm{if $\lambda_{1}-\lambda_{1}^{\prime}\notin{\mathbb{Z}}$};\\ &\sharp\left({\mathrm{A}}_{[d_{1},\cdots,d_{k}]}(\iota)\right)\cdot\sharp\left({\mathrm{A}}_{[d_{1}^{\prime},\cdots,d_{l}^{\prime}]}(\iota^{\prime})\right)\cdot\delta_{\iota,\iota^{\prime}},&\textrm{if $\lambda_{1}-\lambda_{1}^{\prime}\in{\mathbb{Z}}$}\end{aligned}\right.$
@@
    $\delta_{\iota,\iota^{\prime}}=\left\{\begin{aligned} &0,&\textrm{if $\iota\neq\iota^{\prime}$;}\\ &1,&\textrm{if $\iota=\iota^{\prime}$}.\end{aligned}\right.$
@@ 2.1. Associated varieties and genuine representations
Following the notations introduced in Section 1.2 , denote by $({\mathcal{U}}(\mathfrak{g})_{k})_{k=0,1,2,\cdots}$ the natural increasing filtration of ${\mathcal{U}}(\mathfrak{g})$, where ${\mathcal{U}}(\mathfrak{g})_{k}$ is the subspace of ${\mathcal{U}}(\mathfrak{g})$ spanned by elements of the form $X_{1}X_{2}\cdots X_{m}$ for $0\leq m\leq k$, with $X_{j}\in\mathfrak{g}$ for $1\leq j\leq m$. By the Poincaré-Birkhoff-Witt theorem, we can identify the associated ring
@@
is automatically a ${\mathcal{Z}}(\mathfrak{g})$-finite two-sided ideal, we define the complex associated variety $\mathrm{AV}(V)$ of $V$ as the associated variety of the ideal ${\mathrm{Ann}}(V)$, which is a ${\mathrm{Inn}}(\mathfrak{g})$-stable closed subvariety of ${\mathrm{Nil}}(\mathfrak{g}^{*})$. Suppose $S$ is a ${\mathrm{Inn}}(\mathfrak{g})$-stable Zariski closed subset of ${\mathrm{Nil}}(\mathfrak{g}^{*})$, denote by ${\mathrm{Rep}}_{S}(G)$ the category of Casselman-Wallach representations of $G$ whose complex associated variety is contained in $S$.
@@
(2.1) $\psi:$ ${\mathrm{Irr}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q);{\mathcal{O}}\right)$$\longrightarrow$ $\quad{\mathrm{Irr}}_{\nu+\mu}\left({\mathrm{U}}(p,q);{\mathcal{O}}\right)$
@@ 2.2. Generalities on coherent families of Casselman-Wallach representations
It can be verified directly from the above assumption that the homomorphism $\iota$ induces a unique morphism between the abstract Cartan subalgebra of ${\mathrm{Lie}}(G)$ and that of ${\mathrm{Lie}}(G_{{\mathbb{C}}})$. The analytic weight lattice $Q_{an}$ of $G_{\mathbb{C}}$ is identified with a subgroup of ${{}^{a}\mathfrak{h}^{*}}$ via this morphism, and is denoted by $Q_{\iota}\subseteq{{}^{a}\mathfrak{h}^{*}}$. Moreover, $Q_{\iota}$ is $W$-stable and $Q_{\mathfrak{g}}\subseteq Q_{\iota}\subseteq Q^{\mathfrak{g}}$. In the rest of this section, we fix a $Q_{\iota}$-coset $\Lambda=\lambda+Q_{\iota}\subseteq{{}^{a}\mathfrak{h}^{*}}$. Let $W_{\Lambda}$ denote the stabilizer of $\Lambda$ in $W$, namely,
@@ A ${\mathcal{K}}(G)$-valued $\Lambda$-coherent family is a map
We then introduce the notion of parameters for coherent continuation representations and provide an explicit formula for them. Suppose $H$ is a Cartan subgroup of $G$, its Lie complexified Lie algebra is denoted by $\mathfrak{h}$, which is a Cartan subalgebra of $\mathfrak{g}$. $H$ has a unique maximal compact subgroup $T$. Denote by $\Delta_{\mathfrak{h}}\subseteq\mathfrak{h}^{*}$ the root system of $\mathfrak{g}$. A root is called imaginary if $\check{\alpha}\in\mathfrak{t}$, where $\mathfrak{t}=\mathrm{Lie}(T)$, or equivalently if it takes purely imaginary values on $\mathfrak{h}^{*}$. Moreover, an imaginary root $\alpha$ is called compact imaginary if the corresponding root space $\mathfrak{g}_{\alpha}$ is contained in a complexified Lie algebra of a compact subgroup of $G$.
@@
  • Every irreducible Casselman-Wallach representation of $H$ is finite-dimensional.
@@
(2.7) $\Gamma:$ $\Lambda$$\longrightarrow$ $\quad{\mathrm{Irr}}(H)$
@@
where $W(\Delta_{\mathrm{\mathfrak{h},im}})$ is the Weyl group for the imaginary root system $\Delta_{\mathfrak{h},\mathrm{im}}$. There is a group action of $W_{\mathfrak{h},\mathfrak{t}}$ on the set of positive systems of $\Delta_{\mathfrak{h},\mathrm{im}}$, with the subgroup $W(\Delta_{\mathfrak{h},\mathrm{im}})$ acts simply transitively. It is easy to verify that for any choice of positive systems $\Delta_{\mathfrak{h},\mathrm{im}}^{+}$ there is a natural decomposition:
@@
(2.14) $\mathrm{sgn_{\gamma}}:$ $W_{\gamma}$$\longrightarrow$ $\quad{\mathbb{C}}^{\times}$
@@
2.3. Representation theory of Weyl groups
@@
Denote by $\mathrm{S}^{e}(V_{1})$ the $e$-th symmetric power of $V_{1}$. Suppose $\sigma_{1}\in{\mathrm{Irr}}(H_{1})$ is an irreducible representation that occurs with multiplicity $1$ in $\mathrm{S}^{e}(V_{1})$ and does not occur in $\mathrm{S}^{i}(V_{1})$ if $10\leq i\leq e-$ 1. We may regard $\sigma_{1}$ as a subspace of $\mathrm{S}^{e}(V)$ via the inclusion $\sigma_{1}\subseteq{\mathbb{C}}\otimes\mathrm{S}^{e}(V_{1})\subseteq\mathrm{S}^{e}(V)$, and consider the $H$-representation $\sigma$ of $\mathrm{S}^{e}(V)$ generated by $\sigma^{\prime}$. Then
@@
(2.16) $j_{H_{1}}^{H}:$ ${\mathrm{Irr}}(H_{1},e)$$\longrightarrow$ $\quad{\mathrm{Irr}}(H,e)$
@@
Back to our setting of real reductive groups, let $W\subseteq{\mathrm{GL}}({{}^{a}\mathfrak{h}})$ denote the abstract Weyl group. For every $\sigma\in{\mathrm{Irr}}(W)$, its fake degree is defined as
@@
(2.24) ϕ::italic-$\phi:$ $\mathrm{YD}_{n}$$\longrightarrow$ $\quad{\mathrm{Irr}}(\mathrm{S}_{n})$
@@
When we identify nilpotent orbits with partitions of $n$, the bijection $\phi$ coincides with the Springer correspondence between nilpotent orbits and irreducible representations of Weyl groups.